# Make a Cheap Scaling: A Self-Cascade Diffusion Model for Higher-Resolution Adaptation

### v1

@@ Abstract
Diffusion models have proven to be highly effective in image and video generation; however, they still face composition challenges when generating images of varying sizes due to single-scale training data. Adapting large pre-trained diffusion models for higher resolution demands substantial computational and optimization resources, yet achieving a generation capability comparable to low-resolution models remains elusive. This paper proposes a novel self-cascade diffusion model that leverages the rich knowledge gained from a well-trained low-resolution model for rapid adaptation to higher-resolution image and video generation, employing either tuning-free or cheap upsampler tuning paradigms. Integrating a sequence of multi-scale upsampler modules, the self-cascade diffusion model can efficiently adapt to a higher resolution, preserving the original composition and generation capabilities. We further propose a pivot-guided noise re-schedule strategy to speed up the inference process and improve local structural details. Compared to full fine-tuning, our approach achieves a 5×5$\times$5 × training speed-up and requires only an additional 0.002M tuning parameters. Extensive experiments demonstrate that our approach can quickly adapt to higher resolution image and video synthesis by fine-tuning for just $10k$ steps, with virtually no additional inference time. Our code will be released at https://github.com/GuoLanqing/Self-Cascade/.
footnotetext: ${}^{\dagger}$ Equal Contributionsfootnotetext: $*{}^{*}$ Coresponding Authors
@@ 1 Introduction
Figure 1: The average FVD score on the Webvid-10M [1] benchmark for both the full fine-tuning (Full-FT) and our proposed fast adaptation method (Ours) is assessed every $5k$ iterations. We observe that full fine-tuning necessitates a large number of training steps and suffers from poor composition ability and desaturation issues. In contrast, our method enables rapid adaptation to the higher-resolution domain while preserving reliable semantic and local structure generation capabilities.
Over the past two years, stable diffusion (SD) [6, 18] has sparked great interest in generative models, gathering attention from both academic and industry. It has demonstrated impressive outcomes across diverse generative applications, e.g., text-to-image generation [5, 20, 8, 18], image-to-image translation [22, 28], and text-to-video generation [9, 25, 2, 33, 35]. To scale up SD models to high-resolution content generation, a commonly employed approach is progressive training  [6, 18], i.e., training the SD model with lower-resolution images before fine-tuning with higher-resolution images. This warm-up approach enhances the model’s semantic composition ability, leading to the generation of high-quality images. However, even a well-trained diffusion model for low-resolution images demands extensive fine-tuning and computational resources when transferring to a high-resolution domain due to its large size of model parameters. For instance, SD 2.1 [6] requires $550k$ steps of training at resolution $256^{2}$ before fine-tuning with $>\!1000k$ steps at resolution $512^{2}$ to enable $512^{2}$ image synthesis. Insufficient tuning steps may severely degrade the model’s composition ability, resulting in issues such as pattern repetition, desaturation, and unreasonable object structures as shown in Figure 1.
@@
In this paper, we present a novel self-cascade diffusion model that harnesses the rich knowledge gained from a well-trained low-resolution model, to facilitate rapid adaptation to higher resolutions. Our approach begins with the introduction of a tuning-free version, which utilizes a pivot-guided noise re-scheduling strategy. This strategy enforces the synthesis of detailed structures at a new scale by injecting reliable semantic guidance derived from the low-resolution model. Building on this baseline, we further propose time-aware feature upsampling modules for a tuning version to inject the side information from newly acquired higher-resolution data. To enhance the robustness of scale adaptation while preserving the model’s original composition and generation capabilities, we fine-tune a series of plug-and-play, lightweight upsampling modules at different feature levels, using a small amount of acquired high-quality data with a few tuning steps. The proposed upsampler modules can be flexibly plugged into any pre-trained SD-based models, including both image and video generation models. Compared to full fine-tuning, our approach offers a training speed-up of more than 5 times and requires only 0.002M trainable parameters. Extensive experiments demonstrated that our proposed method can rapidly adapt to higher-resolution image and video synthesis with just $10k$ fine-tuning steps and virtually no additional inference time.
@@
  • We propose a novel self-cascade diffusion model by employing a pivot-guided noise re-scheduling strategy to enable tuning-free adaptation to higher resolutions, achieved through the cyclic re-utilization of the low-resolution diffusion model.
  • We further introduce a series of plug-and-play, learnable time-aware feature upsampler modules to incorporate knowledge from a few high-quality images for fine-tuning. This approach achieves a 5×5$\times$5 × training speed-up and requires only an additional 0.002M parameters.
@@ 2 Related Work
Figure 2: Illustration of the proposed self-cascade diffusion model, consisting of (a) Pivot guided noise re-schedule strategy: cyclically re-utilizing the low-resolution model, where the model will progressively adapt to the higher-resolution generation; (b) Time-aware feature upsampler: the denoising process of image $z^{r}_{t}$ in step $t$ will be guided by the pivot guidance $z^{r-1}_{0}$ from the pivot stage (last stage) with a series of plugged-in tuneable upsampler modules.
@@
    Training-free. Recently, Jin et al. [16] explored a training-free approach for variable sizes but did not address high-resolution generation. ScaleCrafter [10] employed dilated convolution to expand the receptive field of convolutional layers for adapting to new resolutions. However, these approaches require careful adjustments, such as dilated stride and injected step, which lack semantic constraints and result in artifacts for various object generation scales.
@@ 3 Preliminary
Our proposed method is based on the recent text-to-image diffusion model (i.e., stable diffusion (SD) [21, 18]), which formulates the diffusion and denoising process in a learned low-dimensional latent space. An autoencoder first conducts perceptual compression to significantly reduce the computational cost, where the encoder $E$ converts image $x_{0}\in\mathbb{R}^{3\times H\times W}$ to its latent code $z_{0}\in\mathbb{R}^{4\times H^{\prime}\times W^{\prime}}$ and the decoder $D$ reconstructs the image $x_{0}$ from the $z_{0}$ as follows,
@@
Then, the diffusion model formulates a fixed forward diffusion process to gradually add noise to the latent code $z_{0}\sim p(x_{0})$:
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})\;.$ (2)
In the inference stage, we sample latent features from the conditional distribution $p(z_{0}|c)$ with the conditional information $c$ (e.g., text embedding with CLIP encoder [19] $E_{CLIP}$):
$p_{\theta}(z_{0:T}|c)=\prod^{T}_{t=1}p_{\theta}(z_{t-1}|z_{t},c).$ (3)
@@
where $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, $\epsilon\in\mathcal{N}(0,\mathbf{I})$ represents the noised feature map at step $t$.
4 Self-Cascade Diffusion Model
4.1 Problem Formulation
@@
Given a pre-trained stable diffusion (SD) model with the denoiser $\epsilon_{\theta}(\cdot)$ for synthesizing low-resolution images (latent code) $z\in\mathbb{R}^{d}$, our goal is to generate higher-resolution images $z^{R}\in\mathbb{R}^{d_{R}}$ in a time/resource and parameter-efficient manner with an adapted model $\tilde{\epsilon}_{\theta}(\cdot)$. To achieve such a goal, we aim to reuse the rich knowledge from the well-trained low-resolution model and only learn the low-level details at a new scale. Thus, we formulate a cascade framework consisting of a low-resolution model at its original scale and auxiliary newly-added modules at the new scale. We intuitively define a scale decomposition to decompose the whole scale adaptation $\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{R}}$ into multiple progressive adaptation processes such that $d=d_{0}<d_{1}\ldots<d_{R}$ where $R=\left\lceil\text{log}_{4}{d_{R}/d}\right\rceil$. We first progressively synthesize a low-resolution image (latent code) $z^{r-1}$ and then utilize it as the pivot guidance to synthesize the higher resolution result $z^{r}$ in the next stage, where the reverse process of cascade diffusion model can be extended by Eq. (3) for each $z^{r}$, $r=1,\ldots,R$ as follows:
$p_{\theta}(z^{r}_{0:T}|c,z^{r-1})=p(z^{r}_{T})\prod^{T}_{t=1}p_{\theta}(z^{r}_{t-1}|z^{r}_{t},c,z^{r-1}),$ (5)
@@
where the reverse transition $p_{\theta}(z^{r}_{t-1}|z^{r}_{t},c,z^{r-1})$ conditions not only on denoising step $t$ and text embedding $c$, but also on lower-resolution latent code $z^{r-1}$ generated in last stage. Previous works, e.g., [12], LAVIE [32], and SHOW-1 [36], approach this problem by adopting extra super-resolution models to perform the conditional image generation where they concat the $z^{r-1}$ and a new initial noise map $z^{r}_{T}$ as the input to the new diffusion denoiser. Different from them, we propose a self-cascade diffusion model to cyclically re-utilize the low-resolution image synthesis model.
4.2 Pivot-Guided Noise Re-Schedule
@@
According to the scale decomposition explained in Section 4.1, the whole scale adaptation process will be decoupled into multiple moderate adaptations, e.g., 4×4$\times$4 × more pixels than the previous stage. The information capacity gap between $z^{r}$ and $z^{r-1}$ is not significant, especially in the presence of noise (intermediate step of diffusion). Consequently, we assume that $p(z_{K}^{r}|z_{0}^{r-1})$ can be considered as the proxy for $p(z_{K}^{r}|z_{0}^{r})$ to manually set the initial diffusion state for current adaptation stage $\mathbb{R}^{d_{r-1}}\rightarrow\mathbb{R}^{d_{r}}$, where $K<T$ is an intermediate step. Specifically, let $\phi_{r}$ denote a deterministic resize interpolation function (i.e., bilinear interpolation) to upsample from scale $d_{r-1}$ to $d_{r}$. We upsample the generated lower-resolution image $z_{0}^{r-1}$ from last stage into $\phi_{r}(z_{0}^{r-1})$ to maintain dimensionality. Then we can diffuse it by $K$ steps to derive $z_{K}^{r-1}$ and use it to replace $z_{K}^{r}$ as follows:
$z_{K}^{r}\sim\mathcal{N}(\sqrt{\bar{\alpha}_{K}}\phi_{r}(z_{0}^{r-1}),\sqrt{1-\bar{\alpha}_{K}}\mathbf{I}).$ (6)
@@
Regarding $z_{K}^{r}$ as the initial state for the current stage and starting denoising with the last $K\rightarrow 0$ steps as Eq. (3) to generate the $z_{0}^{r}$, which is the generated higher-resolution image in the current stage. It is straightforward that such a pivot-guided strategy can be extended to all decoupled scale adaptation stages. Hence, the whole synthesis process for a higher-resolution image with resolution $d_{R}$ using pivot-guided noise re-schedule strategy can be illustrated as:
$z_{T}^{0}\rightarrow\cdots\rightarrow$ $z_{K}^{0}\rightleftharpoons\cdots\rightleftharpoons z_{1}^{0}\rightleftharpoons z_{0}^{0}$
$\downarrow$
$z_{K}^{1}\rightleftharpoons\cdots\rightleftharpoons z_{1}^{1}\rightleftharpoons z_{0}^{1}$
$\ldots$
$\downarrow$
$z_{K}^{R}\rightarrow\cdots\rightarrow z_{1}^{R}\rightarrow z_{0}^{R}.$ (7)
So far, we have devised a tuning-free self-cascade diffusion model to progressively expand the model capacity for higher-resolution adaptation with cyclically re-utilizing the totally frozen low-resolution model as shown in Figure 2(a).
4.3 Time-Aware Feature Upsampler
@@
Although the tuning-free self-cascade diffusion model built upon pivot-guided noise re-schedule (Section 4.2) can achieve a feasible and scale-free higher-resolution adaptation, it has limitations on synthesis performance especially the detailed low-level structures due to the unseen higher-resolution ground-truth images. To achieve a more practical and robust scale adaptation performance, we further introduce a tuning self-cascade diffusion model to enable cheap scaling, by inserting a very lightweight time-aware feature upsampler that can be plugged into any diffusion-based synthesis methods, for more flexible higher-resolution image or video adaptation as illustrated in Figure 2(b).
@@
Specifically, given the denoised image $z^{r}_{t}$ in step $t$ and the pivot guidance $z^{r-1}_{0}$ from the last stage, we can achieve the corresponding intermediate multi-scale feature groups $h_{t}^{r}$ and $h_{0}^{r-1}$ via the pre-trained UNet denoiser $\epsilon_{\theta}$, respectively, as follows:
@@
where $N$ represents the number of features within each feature group. Inspired by the recent work [24] that investigated the impact of various components in the UNet architecture on synthesis performance, we choose to use skip features as a feature group. These features have a negligible effect on the quality of the generated images while still providing semantic guidance. We define a series of time-aware feature upsamplers $\Phi=\{\phi_{1},\phi_{2},\ldots,\phi_{N}\}$ to upsample and transform pivot features at each corresponding scale. During the diffusion generation process, the focus shifts from high-level semantics to low-level detailed structures as the signal-to-noise ratio progressively increases as noise is gradually removed. Consequently, we propose that the learned upsampler transformation should be adaptive to different time steps. The upsampled features $\phi_{n}(h^{r-1}_{n,0},t)$ is then added with original features $h^{r}_{n,t}$ at each scale:
@@
$\mathcal{L}=\mathbb{E}_{z^{r}_{t},z^{r-1}_{0},t,c,\epsilon,t}(\|\epsilon-\tilde{\epsilon}_{\theta+\theta_{\Phi}}(z_{t}^{r},t,c,z^{r-1}_{0})\|^{2}),$ (10)
@@
Extendable tuning. It is noted that our self-cascade diffusion model can be seamlessly extended to a new higher-resolution synthesis, benefiting from the cyclical re-utilization of the shared diffusion model at each scale adaptation stage. For instance, the adapted diffusion model $\tilde{\epsilon}_{\theta}(\cdot)$ for scale $R$ can be further adapted to a previously unseen scale $R^{\prime}$ by re-tuning the integrated feature upsampler modules, without requiring any additional parameters.

### v2

# Project page: https://guolanqing.github.io/Self-Cascade/

@@ Abstract
Diffusion models have proven to be highly effective in image and video generation; however, they encounter challenges in the correct composition of objects when generating images of varying sizes due to single-scale training data. Adapting large pre-trained diffusion models to higher resolution demands substantial computational and optimization resources, yet achieving generation capabilities comparable to low-resolution models remains challenging. This paper proposes a novel self-cascade diffusion model that leverages the knowledge gained from a well-trained low-resolution image/video generation model, enabling rapid adaptation to higher-resolution generation. Building on this, we employ the pivot replacement strategy to facilitate a tuning-free version by progressively leveraging reliable semantic guidance derived from the low-resolution model. We further propose to integrate a sequence of learnable multi-scale upsampler modules for a tuning version capable of efficiently learning structural details at a new scale from a small amount of newly acquired high-resolution training data. Compared to full fine-tuning, our approach achieves a 5×5$\times$5 × training speed-up and requires only 0.002M tuning parameters. Extensive experiments demonstrate that our approach can quickly adapt to higher-resolution image and video synthesis by fine-tuning for just $10k$ steps, with virtually no additional inference time.
Keywords:
Image Synthesis Video Synthesis Diffusion Model Higher-Resolution Adaptation
footnotetext: Equal Contributionsfootnotetext: Corresponding Authors
@@ 1 Introduction
[…]
Over the past two years, stable diffusion (SD) [7, 21] has sparked great interest in generative models, gathering attention from both academic and industry. It has demonstrated impressive outcomes across diverse generative applications, e.g., text-to-image generation [6, 23, 10, 21], image-to-image translation [25, 31], and text-to-video generation [13, 28, 2, 36, 38]. To scale up SD models to high-resolution content generation, a commonly employed approach is progressive training  [7, 21], i.e., training the SD model with lower-resolution images before fine-tuning with higher-resolution images. This warm-up approach enhances the model’s semantic composition ability, leading to the generation of high-quality images. However, even a well-trained diffusion model for low-resolution images demands extensive fine-tuning and computational resources when transferring to a high-resolution domain due to its large size of model parameters. For instance, SD 2.1 [7] requires $550k$ steps of training at resolution $256^{2}$ before fine-tuning with $>\!1000k$ steps at resolution $512^{2}$ to enable $512^{2}$ image synthesis. Insufficient tuning steps may severely degrade the model’s composition ability, resulting in issues such as pattern repetition, desaturation, and unreasonable object structures as in Fig. 1.
@@
In this paper, we present a novel self-cascade diffusion model that harnesses the rich knowledge gained from a well-trained low-resolution generation model to enable rapid adaptation to higher-resolution generation. Our approach begins with the introduction of a tuning-free version, which utilizes a pivot replacement strategy to enforce the synthesis of detailed structures at a new scale by injecting reliable semantic guidance derived from the low-resolution model. Building on this baseline, we further propose time-aware feature upsampling modules as plugins to a base low-resolution model to conduct a tuning version. To enhance the robustness of scale adaptation while preserving the model’s original composition and generation capabilities, we fine-tune the plug-and-play and lightweight upsampling modules at different feature levels, using a small amount of acquired high-quality data with a few tuning steps.
The proposed upsampler modules can be flexibly plugged into any pre-trained SD-based models, including both image and video generation models. Compared to full fine-tuning, our approach offers a training speed-up of more than 5 times and requires only 0.002M trainable parameters. Extensive experiments demonstrated that our proposed method can rapidly adapt to higher-resolution image and video synthesis with just $10k$ fine-tuning steps and virtually no additional inference time.
@@
  • We propose a novel self-cascade diffusion model for fast-scale adaptation to higher resolution generation, by cyclically re-utilizing the low-resolution diffusion model. Based on that, we employ a pivot replacement strategy to enable a tuning-free version as the baseline.
  • We further construct a series of plug-and-play, learnable time-aware feature upsampler modules to incorporate knowledge from a few high-quality images for fine-tuning. This approach achieves a 5×5$\times$5 × training speed-up compared to full fine-tuning and requires only 0.002M learnable parameters.
@@ 2 Related Work
    Tuning-free. Several methods [19, 12, 8, 11] have explored expanding low-resolution diffusion models to higher resolutions without tuning. Recently, Jin et al. [19] explored a tuning-free approach for variable sizes but did not address high-resolution generation. ScaleCrafter [12] employed dilated convolution to expand the receptive field of convolutional layers for adapting to new resolutions. Besides, DemoFusion [8] used low-resolution semantic guidance as well as dilated sampling strategy to achieve a high-resolution generation. However, these approaches require careful adjustments, such as dilated stride and injected step, which lack semantic constraints and result in artifacts for various generation scales.
@@ 3 Preliminary
Our proposed method is based on the recent text-to-image diffusion model (i.e., stable diffusion (SD) [24, 21]), which formulates the diffusion and denoising process in a learned low-dimensional latent space. An autoencoder first conducts perceptual compression to significantly reduce the computational cost, where the encoder $E$ converts image $x_{0}\in\mathbb{R}^{3\times H\times W}$ to its latent code $z_{0}\in\mathbb{R}^{4\times H^{\prime}\times W^{\prime}}$ and the decoder $D$ reconstructs the image $x_{0}$ from $z_{0}$ as follows,
@@
Then, the diffusion model formulates a fixed forward diffusion process to gradually add noise to the latent code $z_{0}\sim p(z_{0})$:
$q(z_{t}|z_{0})=\mathcal{N}(z_{t};\sqrt{\bar{\alpha}_{t}}z_{0},(1-\bar{\alpha}_{t})\mathbf{I})\;.$ (2)
In the inference stage, we sample latent features from the conditional distribution $p(z_{0}|c)$ with the conditional information $c$ (e.g., the text embedding extracted by a CLIP encoder [22] $E_{CLIP}$):
$p_{\theta}(z_{0:T}|c)=p(z_{T})\prod^{T}_{t=1}p_{\theta}(z_{t-1}|z_{t},c).$ (3)
@@
where $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ represents the noised feature map at step $t$.
Figure 2: Illustration of the proposed self-cascade diffusion model, which is implemented in both tuning-free and tuning versions. (a) For the tuning-free version, we cyclically re-utilize the low-resolution model to progressively adapt it to the higher-resolution generation; (b) For the tuning version, we additionally plug feature upsamplers ($\Phi$) into the base low-resolution generation model: the denoising process of image $z^{r}_{t}$ in step $t$ will be guided by the pivot guidance $z^{r-1}_{0}$ from the pivot stage (last stage) with a series of plugged-in tuneable upsampler modules.
4 Methodology
In this section, we first introduce the overall framework of the proposed self-cascade diffusion model (Sec. 4.1). Based on that, we propose a tuning-free version using a pivot replacement strategy as the baseline, as well as an improved tuning version by plugging tunable feature upsamplers (Sec. 4.2). We then provide an analysis and discussion on our self-cascade diffusion model (Sec. 4.3).
4.1 Self-Cascade Diffusion Model
@@
Given a pre-trained stable diffusion (SD) model with the denoiser $\epsilon_{\theta}(\cdot)$ for synthesizing low-resolution images (latent code) $z\in\mathbb{R}^{d}$, our goal is to generate higher-resolution images $z^{R}\in\mathbb{R}^{d_{R}}$ in a time/resource and parameter-efficient manner with an adapted model $\tilde{\epsilon}_{\theta}(\cdot)$. To achieve such a goal, we aim to reuse the rich knowledge from the well-trained low-resolution model and only learn the low-level details at a new scale. Thus, we propose a self-cascade diffusion model to cyclically re-utilize the low-resolution image synthesis model. We intuitively define a scale decomposition to decompose the whole scale adaptation $\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{R}}$ into multiple progressive adaptation processes such that $d=d_{0}<d_{1}\ldots<d_{R}$ where $R=\left\lceil\text{log}_{4}{d_{R}/d}\right\rceil$. We first progressively synthesize a low-resolution image (latent code) $z^{r-1}$ and then utilize it as the pivot guidance to synthesize the higher resolution result $z^{r}$ in the next stage, where the reverse process of the self-cascade diffusion model can be extended by Eq. (3) for each $z^{r}$, $r=1,\ldots,R$ as follows:
$p_{\theta}(z^{r}_{0:T}|c,z^{r-1}_{0})=p(z^{r}_{T})\prod^{T}_{t=1}p_{\theta}(z^{r}_{t-1}|z^{r}_{t},c,z^{r-1}_{0}),$ (5)
@@
where the reverse transition $p_{\theta}(z^{r}_{t-1}|z^{r}_{t},c,z^{r-1}_{0})$ not only conditions on denoising step $t$ and text embedding $c$, but also on lower-resolution latent code $z^{r-1}_{0}$ generated in the last stage. Different from previous works, e.g., [16], LAVIE [35], and SHOW-1 [39] where $p_{\theta}$ in Eq. 9 is implemented by a new super-resolution model, we cyclically re-utilize the base low-resolution synthesis model to inherit the prior knowledge of the base model thus improve the efficiency.
@@
Pivot replacement. According to the scale decomposition, the whole scale adaptation process will be decoupled into multiple moderate adaptation stages, e.g., 4×4$\times$4 × more pixels than the previous stage. The information capacity gap between $z^{r}$ and $z^{r-1}$ is not significant, especially in the presence of noise (intermediate step of diffusion). Consequently, we assume that $p(z_{K}^{r}|z_{0}^{r-1})$ can be considered as the proxy for $p(z_{K}^{r}|z_{0}^{r})$ to manually set the initial diffusion state for current adaptation stage $\mathbb{R}^{d_{r-1}}\rightarrow\mathbb{R}^{d_{r}}$, where $K<T$ is an intermediate step. Specifically, let $\phi$ denote a deterministic resize interpolation function (i.e., bilinear interpolation) to upsample from scale $d_{r-1}$ to $d_{r}$. We upsample the generated lower-resolution image $z_{0}^{r-1}$ from last stage into $\phi(z_{0}^{r-1})$ to maintain dimensionality. Then we can diffuse it by $K$ steps and use it to replace $z_{K}^{r}$ as follows:
$z_{K}^{r}\sim\mathcal{N}(\sqrt{\bar{\alpha}_{K}}\phi(z_{0}^{r-1}),\sqrt{1-\bar{\alpha}_{K}}\mathbf{I}).$ (6)
@@
Regarding $z_{K}^{r}$ as the initial state for the current stage and conduct denoising with $K\rightarrow 0$ steps as Eq. (3) to generate the $z_{0}^{r}$, which is the generated higher-resolution image in the current stage.
@@
We can employ such a pivot replacement strategy at all decoupled scale adaptation stages. Hence, the whole synthesis process for a higher-resolution image with resolution $d_{R}$ using pivot replacement strategy can be illustrated as Fig. 2(a). So far, we have devised a tuning-free version of self-cascade diffusion model (denoted as Ours-TF in experiments) to progressively expand the model capacity for higher-resolution adaptation with cyclically re-utilizing the totally frozen low-resolution model. Although the tuning-free self-cascade diffusion model built upon pivot replacement strategy (Sec. 4.1) can achieve a feasible and scale-free higher-resolution adaptation, it has limitations on synthesis performance especially the detailed low-level structures due to the unseen higher-resolution ground-truth images. To achieve a more practical and robust scale adaptation performance, we further introduce an improved tuning version of the self-cascade diffusion model (denoted as Ours-T in experiments) in Sec. 4.2.
4.2 Feature Upsampler Tuning
In this section, we propose a tuning version of the self-cascade diffusion model that enables a cheap scaling, by inserting very lightweight time-aware feature upsamplers as illustrated in Fig. 2(b). The proposed upsamplers can be plugged into any diffusion-based synthesis methods. The detailed tuning and inference processes of our tuning version self-cascade diffusion model are in Algorithm 1 and 2, respectively. Note that by omitting the tuning process and solely executing the inference step in Algorithm 2, it turns into our tuning-free version.
@@ 4.2 Feature Upsampler Tuning
Specifically, given an intermediate $z^{r}_{t}$ in step $t$ and the pivot guidance $z^{r-1}_{0}$ from the last stage, we can achieve corresponding intermediate multi-scale feature groups $h_{t}^{r}$ and $h_{0}^{r-1}$ via the pre-trained UNet denoiser $\epsilon_{\theta}$, respectively, as follows:
@@
where $N$ represents the number of features within each feature group and the details are included in the supplementary. In short, inspired by the recent work [27] that investigated the impact of various components in the UNet architecture on synthesis performance, we choose to use skip features as a feature group. These features have a negligible effect on the quality of the generated images while still providing semantic guidance. We define a series of time-aware feature upsamplers $\Phi=\{\phi_{1},\phi_{2},\ldots,\phi_{N}\}$ to upsample and transform pivot features at each corresponding scale. During the diffusion generation process, the focus shifts from high-level semantics to low-level detailed structures as the signal-to-noise ratio progressively increases as noise is gradually removed. Consequently, we propose that the learned upsampler transformation should be adaptive to different time steps. The upsampled features $\phi_{n}(h^{r-1}_{n,0},t)$ is then added to original features $h^{r}_{n,t}$ at each scale:
@@
$\mathcal{L}=\mathbb{E}_{z^{r}_{t},z^{r-1}_{0},t,c,\epsilon,t}(\|\epsilon-\tilde{\epsilon}_{\theta,\theta_{\Phi}}(z_{t}^{r},t,c,z^{r-1}_{0})\|^{2}),$ (9)